# From Captions to Keyframes: KeyScore for Multimodal Frame Scoring and Video-Language Understanding

### v1

@@ Abstract
Efficient video–language understanding requires selecting a small set of frames that retain semantic and contextual information from long videos. We propose KeyScore, a multimodal frame scoring framework that jointly leverages captions and visual context to estimate frame-level importance. By combining semantic similarity, temporal diversity, and contextual drop impact, KeyScore identifies the most informative frames for downstream tasks such as retrieval, captioning, and video–language reasoning. To complement KeyScore, we introduce STACFP (Spatio-Temporal Adaptive Clustering for Frame Proposals), which generates compact and diverse frame candidates for long-form videos. Together, these modules achieve up to 99% frame reduction compared to full-frame inference and substantially outperform standard 8-frame encoders on MSRVTT, MSVD, and DiDeMo. Our results demonstrate that emphasizing multimodal alignment between visual and textual signals enables scalable, efficient, and caption-grounded video understanding — without explicit video summarization.
@@ 4.2 Keyframe Extraction
Method F1 Method F1

### v2

@@ Abstract
Selecting informative keyframes is critical for efficient video understanding, yet existing approaches often rely on heuristics, ignore semantics, or produce redundant frames. We propose KeyScore, a caption-aware frame scoring method that combines three complementary signals: semantic similarity to captions, temporal representativeness, and contextual drop impact. Applied to large-scale video–caption datasets, KeyScore generates frame-level importance scores that enable training keyframe extractors or guiding video–language models. To support this, we also propose STACFP—a Spatio-Temporal Adaptive Clustering method that generates diverse and compact frame proposals across long videos. Together, KeyScore and STACFP reduce uninformative frames while preserving critical content, resulting in faster and more accurate inference. Our experiments on three standard video-language benchmarks (MSRVTT, MSVD, DiDeMo) show that combining STACFP and KeyScore enables up to 99% frame reduction compared to full-frame processing, while outperforming uniform 8-frame encoders in video-text retrieval, keyframe extraction, and action recognition tasks. By focusing on semantically relevant frames, our method enhances both efficiency and performance, enabling scalable and caption-grounded video understanding.
@@ 4.2 Keyframe Extraction
Method F1-Score Method F1-Score